OpenStudy (girlgoyleh.):

Which set of integers is a Pythagorean triple? A. 15, 9, 17 B. 6, 8, 12 C. 3, 5, 6 D. 5, 12, 13

OpenStudy (anonymous):

D is one

OpenStudy (girlgoyleh.):

i thought it would be A.

OpenStudy (anonymous):

13^2=169 12^2 + 5^2 =144+25=169

OpenStudy (anonymous):

no its not D

OpenStudy (anonymous):

its C.

OpenStudy (anonymous):

i just did the math its C.

OpenStudy (girlgoyleh.):

I think it is D, try using a calculator to check your work.

OpenStudy (anonymous):

no its C

OpenStudy (anonymous):

A is not \[ 17^2=289\\ 9^2+15^2=306 \]

OpenStudy (anonymous):

i had the same qestion and i chose d and got it right

OpenStudy (girlgoyleh.):

then why is it c if you got d right

OpenStudy (anonymous):

i accidentaly typed d when he said d

OpenStudy (anonymous):

i was reading his commnet when i was typing

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

C is not \[ 6^2=36\ne 3^2 + 5^2=34 \]

OpenStudy (girlgoyleh.):

could you help with another one @eliassaab

OpenStudy (anonymous):

Post it as a new problem and tag me in it